OpenStudy (anonymous):

Fitzgerald purchased a car for $18,980. He made a down payment of $2,240. He applied for a five-year installment loan with an interest rate of 7.4%. What is the total cost of the car after five years?

OpenStudy (anonymous):

$20,078.40 $22,318.40 $17,978.76

OpenStudy (tkhunny):

Payments made annually?

OpenStudy (tkhunny):

It won't be less than $18,980, so you can just throw that bottom one out!

OpenStudy (anonymous):

Doesn't say it's annually or not so I'm not sure

OpenStudy (anonymous):

I was thinking $22,318.40

OpenStudy (tkhunny):

Well, that's just silly. The problem should say what it wants. Let's do five annual payments at the end of the year. 18980 - 2240 = 16740 -- This is the amount financed. I = 0.074 -- Annual interest rate v = 1/(1+i) Pmt we don't know. We'll have to calculate it. 16740 = Pmt(v + v^2 + v^3 + v^4 + v^5) Do you believe this?

OpenStudy (anonymous):

What I did was 16740(1+0.074)^5 and with that I got $23,920.86 and that answer was close to $22,318.40 so I was thinking that's the answer

OpenStudy (tkhunny):

That's no good. There is only one payment. If you make no payment at all, that is the right answer. This Fitzgerald is making payments along the way. I should have said: v = 1/(1+i) -- Annual Discount Factor Okay, NOW do you believe my expression?

OpenStudy (anonymous):

What do you mean do I believe it?

OpenStudy (tkhunny):

You need to buy into this expression. 16740 = Pmt(v + v^2 + v^3 + v^4 + v^5)

OpenStudy (anonymous):

Okay? I've never been taught that formula though so I'm pretty confused

OpenStudy (whpalmer4):

oh, the loan is almost certainly monthly payments!

OpenStudy (tkhunny):

It's not a formula. It's basic principles. You may have a formula for the present value of five payments at the end of each period. I'm just building it from scratch.

OpenStudy (whpalmer4):

and indeed, the answer for monthly payments is among the choices.

OpenStudy (anonymous):

So is it $22,318.40? Like I said?

OpenStudy (tkhunny):

No, your version still had only one payment at the end of the 5 years. It should be nowhere near that.

OpenStudy (anonymous):

Then $20,078.40?

OpenStudy (tkhunny):

Are you happy with guessing and narrowing or would you like to prove it?

OpenStudy (anonymous):

I don't know how to prove it that's the thing and your "basic principal" theory really confuses me... So if you want to explain that....

OpenStudy (tkhunny):

Do you have a formula for the present value of some number of payments at the end of each period?

OpenStudy (tkhunny):

You either have a formula or you have to build it. Those are the only two choices.

OpenStudy (anonymous):

B = P(1 + i)nt

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

or M = B i(1+i)^nt / (1+i)nt - 1

OpenStudy (tkhunny):

Not enough payments. \(B = \dfrac{1 - v^{n}}{i}\), maybe?

OpenStudy (tkhunny):

Not sure. What are M and B?

OpenStudy (anonymous):

B is 18980, M is just what the monthly payment will equal once you plug everything in.

OpenStudy (anonymous):

May I ask, what is this "v" you keep referring to?

OpenStudy (tkhunny):

That's why I defined it. It was the ANNUAL discount factor. It's how these formulas are built. Now that we know it's monthly, thanks to @whpalmer4's observation, we'll need a new one. I = 0.074 -- Annual interest rate j = I/12 -- Monthly interest rate v = 1/(1+j) -- Monthly discount factor. It works like this. Right now, a payment right now is worth Pmt Right now, a payment one month from now is worth Pmt*v Right now, a payment two months from now is worth Pmt*v^2 Right now, a payment three months from now is worth Pmt*v^3 Making any sense?

OpenStudy (anonymous):

No

OpenStudy (tkhunny):

If we can't build the formula, you must find the right one. That's all there is to it. I cannot help you with that because I can't see what formulas you have in front of you.